please help me to simplify the expression ((x^2-16)/2x)(12/(6x-24)).
Thank You

Answer by Maths68(1474)

=((x^2-16)/2x)(12/(6x-24))
Let A=x^2-16/2x
=x^2-4^2/2x
=(x-4) (x+4)/2x

Let B=12/(6x-24)
=12/6(x-4)
=2/(x-4)
=A*B
=[(x-4) (x+4)/2x]* [2/(x-4)]
=(x+4)/x